A weak base with a Kb value of 4 x 10-8. The pH of a 0.01 solution of the weak base is in the range of


  • 7 to 8
  • 4 to 5
  • 9 to 10
  • 11 to 12

(Explain why)



1
Expert's answer
2021-04-25T23:15:13-0400

Range of 11 to 12.


The value of K_b


Kb

​ for caffine = "4\times 10^{-4}"


4×10−4

CafOH(aq)leftharpoons "Caf(aq)+OH^-(aq)"


CafOH(aq)leftharpoonsCaf(aq)+OH

(aq)

Initial

  0     0.01 M   0

AT equilibrium:

 x     (0.01 -x)M   x

"K_b=\frac{x(0.01-x)}{(x)}"


"Kb\n\n\u200b=(x)\n\n\n\nx(0.01\u2212x)\n\n\u200b\n\n4\\times 10^{-4}=\\frac{x(0.01-x)}{(x)}\n\n\n\n4\u00d710\u22124\n\n=(x)\n\n\n\nx(0.01\u2212x)"

Solving for x:

x = 0.0096 M

The pOH of the solution is given by :

"pOH=-\log[{OH^-}]"

pOH=-\log[x]


pOH=−log[x]

pOH="-\log[0.0096]"


pOH="-" log[0.0096]

pOH = 2.02

pH= 14 "-" pOH = 14 "-" 2.02 = 11.98